OpenStudy (anonymous):

1/cot^2theta+sec theta cos theta

OpenStudy (anonymous):

what do you need?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

\[\frac{ 1 }{ \cot ^2 \theta }+\sec \theta \cos \theta\] is this what you have?

OpenStudy (anonymous):

need to find expression as a constant a single trigo function

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so \[\frac{ 1 }{\cot \theta }=\tan \theta\] and \[\sec \theta = \frac{ 1 }{ \cos \theta }\] what does that get you?

OpenStudy (anonymous):

how did u got this stuff

OpenStudy (anonymous):

that's just the definition

OpenStudy (anonymous):

i really dont get this question at all can u help me plz

OpenStudy (anonymous):

\[\tan \theta = \frac{ \text{opposite} }{ \text{adjacent} }\text{, }\cot \theta = \frac{ \text{adjacent} }{ \text{opposite} }\Rightarrow \frac{ 1 }{ \cot \theta } = \tan \theta\]\[\sec \theta = \frac{ \text{hypoteneuse} }{ \text{opposite}}\text{, }\cos \theta = \frac{ \text{opposite} }{ \text{hypoteneuse}}\Rightarrow \sec \theta = \frac{1}{\cos \theta}\]

OpenStudy (anonymous):

so substitute... \[\frac{ 1 }{ \cot ^2 \theta }= \tan^2 \theta\] and \[\sec \theta \cos \theta = \frac{ 1 }{ \cos \theta }\cos \theta = 1\] then you have the identity that \[\tan^2\theta +1 = \sec^2 \theta\]

OpenStudy (anonymous):

so thats the answer

OpenStudy (anonymous):

hey u their

OpenStudy (anonymous):

their?

OpenStudy (anonymous):

tan2 theta +1=sec^2 theta is the answer right

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

thanks buddy

OpenStudy (anonymous):

you're welcome